# Boosting Performance on ARC is a Matter of Perspective

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated extraordinary capabilities across diverse tasks, from natural language processing to code generation. Even so, evaluating the extent to which these systems possess abstract reasoning abilities continues to pose a major challenge in the artificial intelligence community. The Abstraction and Reasoning Corpus (ARC-AGI), introduced by Chollet (2019) and designed to assess core knowledge and the ability to generalize in AI, exemplifies this difficulty. Although these tasks (as illustrated in Figure 1) may appear straightforward to humans, both traditional algorithmic approaches (Wind, 2020) and contemporary neural architectures (Li et al., 2024a) have struggled to achieve significant success on ARC-AGI, highlighting potential limitations in current artificial reasoning methods.
Figure 1: Example of a typical ARC-AGI tasks.
Although scaling up models has undoubtedly yielded substantial performance gains on many tasks, size alone does not fully address the core limitations evident in challenges like ARC-AGI. Indeed, the rapid evolution of open-source systems – such as LLaMA-3.2-3B (Dubey et al., 2024) and Nvidia NeMo-Minitron-8B (Sreenivas et al., 2024) – demonstrates that significant capabilities can emerge even at more modest scales. This aligns with mounting evidence that many perceived shortcomings in large language models stem from implementation details or suboptimal data representations rather than from fundamental reasoning deficits (Singh & Strouse, 2024b; Bostrom & Durrett, 2020; Sun et al., 2023). For instance, Allen-Zhu & Li (2024) observe that models may be aware of their mistakes without being able to correct them, while Allen-Zhu & Li (2023) highlight how subtle data modeling choices can impede fine-tuning progress. Collectively, these insights suggest that models often possess the latent capacities needed to tackle ARC-AGI; the real challenge is creating the conditions under which these capacities can be reliably expressed.
@@
In the sections that follow, we detail our data modeling and training strategies, describe our DFS-based solution exploration, and provide comprehensive results and ablation studies. Upon acceptance, we will make our final model, along with the training and inference code, publicly available.
@@
Ryan Greenblatt (Greenblatt, ) 42
Jeremy Berman (Berman, ) 58.5
@@ 2 Related Work
The Abstraction and Reasoning Corpus (ARC) has played a central role in advancing research on abstract reasoning in artificial intelligence, inspiring a wide range of studies focused on its dataset, competitive benchmarks, and the development of innovative solutions driven by resource constraints.
@@
The Abstraction and Reasoning Corpus (ARC-AGI) introduced by Chollet (2019) challenges the idea that language models can effectively generalize from a small number of examples, often referred to as few-shot prompting. The original ARC-AGI dataset consists of 900 reasoning tasks, divided into 400 training tasks, 400 public evaluation tasks, and 100 private and thus unpublished evaluation tasks. Each task involves grids of varying sizes, ranging from 1x1 to 30x30 and utilize a palette of ten distinct colors.
The objective of the individual ARC tasks is to discern this rule from the examples and apply it to new input grids to generate the correct outputs. A task is considered successfully solved when the model produces the accurate output within a maximum of two attempts. Designed to be straightforward for humans yet challenging for machine learning systems, the tasks highlight the current limitations of AI in abstract reasoning. In a study by LeGris et al. (2024), the average human was able to correctly solve 60.2% of the evaluation tasks, while 97.8% of the tasks were solved by at least one participant using two guesses.
Visual Representation of a Task Instance:
Input                                              Output   2 1 1 0 1 2 1          1 1 1
1 2 1 0 2 2 2          1 3 1
2 1 1 0 1 1 1          1 1 1
Compact String Format of same Instance:
<bos>A
Z az I 2 1 1 0 1 2 1 LF
1 2 1 0 2 2 2 LF
2 1 1 0 1 1 1 LF
O 1 1 1 LF
1 3 1 LF
1 1 1 LF
<eos>
Figure 2: Our standard tokenization approach. Note that we use one token per cell instead of compressing the problem more. We also try to not include any unnecessary delimiters. The Pre-prompt (“A…Za…z”) is only included for the first example. Depending on the model and run there might be some small changes to the pre-prompt and input/output prefix tokens.
@@
Other approaches explored various strategies for utilizing LLMs. In Li et al. (2024c), the authors classify two different avenues: Induction, where a LLM infers a function that can solve the problem which is then applied (often using python or a DSL), and Transduction, where the LLM directly generates the solution using a tokenized description of the problem (see Figure 2). The authors argue that these approaches solve different kinds of problems, despite using the same underlying architecture. In their experiments, induction and transduction solve roughly the same amount of problems ($38\%$ and $43\%$ respectively), which can be increased to $56.75\%$ by employing ensembles. Additionally, they use the induction network to generate a large set of novel challenges, dubbed ARC-Heavy. Additionally, some approaches make use of alternative ARC datasets, such as ConceptARC (Moskvichev et al., 2023). The most notable - and only additional dataset we use - is the well-known RE-ARC dataset. Hodel (2024) introduced this dataset, which re-implements all $400$ tasks of the public training dataset. Their code can be used to generate an arbitrary amount of training data for these tasks, but does not introduce novel challenges. All other datasets might include challenges that mimic the evaluation challenges - thereby reducing the difficulty of those challenges immensely. By only using the RE-ARC dataset we still increase our training data immensely, but stay close to solving the ARC challenge as intended.
Data augmentation has been a common approach in previous ARC-AGI competitions (Akyürek et al., 2024; Li et al., 2024b). However, our method extends beyond traditional dataset augmentation, applying transformations throughout our approach, during training, inference and selection.
@@ 3 Notations and Setup
over candidate solutions $s\in\mathcal{S}_{p}$. If exactly one possible valid answer exists the distribution $P(\cdot\mid p)$ would be sharply peaked at $s^{*}_{p}$. While this is the case for most challenges, we will keep our theory more general, assuming multiple valid answers might exist. This can also arise from insufficient information in the given example pairs, which in the worst case prevents us from uniquely inferring the correct solution based solely on the provided data. Examples for this are sometimes found in ARC-AGI, which frequently results in an update of the dataset (Neoneye, ; RubenKelevra, ).
@@
Figure 4: Accuracy and coverage of different selection methods as a function of the confidence threshold $T$. The solid black line shows the proportion of tasks where the correct solution is among the generated candidates. The solid colored lines show what percentage of the tasks would be solved using different aggregation methods (top-2 accuracy), while the dotted lines show how this percentage relates to the black line. Our product of probabilities approach performs best among the tested aggregation methods.
@@ 4 Methods
The trained language model (LLM) then provides us with a probability distribution $\hat{P}(\cdot\mid p)$ over solutions $\mathcal{S}_{p}$. Using multinomial sampling, this allows us to sample $s\sim\hat{P}(\cdot\mid p)$ However, sampling repeatedly from $\hat{P}$ may be expensive and does not ensure coverage of high-probability solutions. In contrast, enumerating $\mathcal{S}_{p}$ in full would provide us with full knowledge of $\hat{P}(s)$ but is impractical. Instead, we rely on a more systematic procedure to select promising candidates by deriving a candidate set of high-probability solutions via threshold-based search. Subsequently, we refine their probabilities by aggregating over multiple problem augmentations.
@@
However, once we have generated the set $\mathcal{C}_{p,T}$, the highest-probability solution according to a single augmentation is not always correct, partly due to autoregressive inconsistencies.
@@
This two-step approach — (1) DFS-based generation with single-augmentation pruning and (2) post-hoc multi-augmentation scoring — ensures that we systematically explore high-probability solutions and then refine their rankings, accounting for LLM inconsistencies across problem representations. In practice, even if several solutions enter $\mathcal{C}_{p,T}$, their final ranks can vary greatly. By consolidating evidence from multiple perspectives, the correct solution often stands out and becomes easier to pinpoint.
@@ 4.1 Product of Expert Augmentations
See Appendix A for a proof of Theorem 4.1. The key takeaway is that $\log Z\leq 0$ becomes smaller whenever augmentations disagree, which can improve the ensemble in expectation relative to any random single-augmentation predictor. As a result, this approach performs especially well when different experts disagree - a state which naturally arises in our case, due to the causal autoregressive nature of the LLMs.
@@
Table 2: Two-guess-accuracy on the ARC-AGI public evaluation set when adding parts of our method. Baseline score shows performance of our network after initial fine-tuning, generating two samples with stochastic sampling. TTT adds test-time training. 16xAug uses augmentation to generate a candidate for each of 16 different augmentations of each task, choosing the two with highest sampling probability as guesses. PoE uses the product of expert augmentations approach with 16 augmentations. Finally, DFS leverages our custom depth-first-search sampling scheme with $T=9\%$ for each of the 16 augmentations.
@@ 5.1 Data Modeling
[…]
@@
As illustrated in Figure 2, we add a small number of extra tokens to the start of a task. Surprisingly, this addition improves the model’s performance. We believe that during fine-tuning (where the embedding layers are also trained), the model learns to use these extra tokens as a form of computational buffer, which influences every subsequent token, thereby enhancing overall performance.
@@ 5.2 Training the models
Therefore, we used Low-Rank Adaptation (LoRA)(Hu et al., 2021), 4-bit quantization and gradient checkpointing, all supported by the unsloth library. We applied the LoRA adaptations to all layers of the network, including the input and output embeddings.
@@
Initial fine-tuning: The initial fine-tuning used a LoRA rank of $256$ and was done on a single H100 GPU. While several ARC-like datasets exist, such as Concept-ARC (Moskvichev et al., 2023), and ARC-Heavy (Li et al., 2024b), we elect to only use RE-ARC (Hodel, 2024) for training. This is done to minimize ”conceptual leakage”, where a particular type of problem might be present in the training data, reducing the difficulty of the evaluation tasks substantially in a way that was not intended. Instead, we train only on replications of the training examples of the offical ARC-AGI training set (i.e. RE-ARC), minimizing this effect and making sure that our results are robust.
@@
The initial fine-tuning took 98 GPU hours on a Nvidia H100 while test-time training takes on average 51 seconds for a single task measured on a Nvidia RTX 4090 GPU. For an overview of our training parameters, see Table 4 in the Appendix.
@@ 5.3 Solution Inference
As introduced in Section 4, we generate candidates for potential solutions using DFS-based sampling to produce the set $C_{p,T}$. The goal here is to generate a small set of candidates with a high chance of containing the correct solution - and doing so quickly. Our set of augmentations $\Phi$ includes $16$ functions per task - each $D_{8}$ symmetry is used twice but with different, randomly chosen color permutations and example re-orderings. Table 3 provides a comparison between DFS, Beam-search, and multinomial and greedy sampling. DFS sampling is able to quickly and efficiently find a high quality set of candidates, while having low computational overhead compared to stochastic sampling for generating multiple solutions and using substantially less VRAM than beam search. In addition, it exhibits a lower false positive rate. While DFS with $T=9\%$ finds less correct solutions than 4x Stochastic sampling ($76.0\%$ vs $77.3\%$), it still results in a better selection score, as it, on average, only returns about half as many false positives. Moreover, DFS accomplished this using only a fourth of the inference time (9:32h vs 39:47h).
While beam search with 4 beams can achieve the same performance as DFS with $T=9\%$, it requires roughly twice the amount of VRAM (7.3GB vs 14GB) and 4 times as much time (37:36h vs 9:32h). It is important to note that comparisons with beam search should be interpreted with caution, as the beam search algorithm is not implemented in the unsloth library used for the other experiments, which might provide some time savings. However, even when accounting for those savings, beam search still requires far more time overall, as it returns a significant amount of false positives, which increase the runtime required in the subsequent selection process, where each candidate is evaluated under different augmentations.
We compare the number of candidate sets that contain the correct solution for different values of $T$ in Figure 4. This function is monotonically increasing in $T$, but so are inference costs and the size of the set $C_{p,T}$, making a selection of the correct candidate harder. Our final results are calculated using $T=9\%$, as it uses roughly the same amount of inference time as greedy sampling.
We provide pseudo-code for the DFS sampling algorithm in Algorithm 1 in the Appendix.
Table 3: Comparison of different sampling strategies on the 400 tasks of the ARC-AGI public evaluation set: Under “Candidate generation”, we list the percentage of correct solutions sampled with different strategies using 16 augmented versions (reflections and rotations, as well as randomly permuted colors and examples) of each task. We also list the average number of candidates generated per task, the runtime of the sampling process on the full dataset and the maximum video memory consumption. Under “Selection”, we compare the accuracy of various selections strategies, performed on the scores calculated in a subsequent scoring process on 16 additional random augmentations. Total runtime includes the test-time fine-tuning on a task’s examples (see Table 4 in the appendix), as well as the candidate generation and selection process. All experiments were performed with the NeMo-Minitron-8B model on a Nvidia RTX 4090 GPU.

### v2

# Product of Experts with LLMs: Boosting Performance on ARC Is a Matter of Perspective

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated extraordinary capabilities across diverse tasks, from natural language processing to code generation. Even so, evaluating the extent to which these systems possess abstract reasoning abilities continues to pose a major challenge in the artificial intelligence community. The Abstraction and Reasoning Corpus (ARC-AGI), introduced by Chollet (2019) and designed to assess core knowledge and the ability to generalize in AI, exemplifies this difficulty. Although these tasks (as illustrated in Figure 1) may appear straightforward to humans, both traditional algorithmic approaches (Wind, 2020) and contemporary neural architectures (Li et al., 2024) have struggled to achieve significant success on ARC-AGI, highlighting potential limitations in current artificial reasoning methods.
Figure 1: Example of a typical ARC-AGI task.
Although scaling up models has undoubtedly yielded substantial performance gains on many tasks, size alone does not fully address the core limitations evident in challenges like ARC-AGI. Indeed, the rapid evolution of open-source systems – such as LLaMA-3.2-3B (Dubey et al., 2024) and Nvidia NeMo-Minitron-8B (Sreenivas et al., 2024) – demonstrates that significant capabilities can emerge even at more modest scales. This aligns with mounting evidence that many perceived shortcomings in large language models stem from implementation details or suboptimal data representations rather than from fundamental reasoning deficits (Singh & Strouse, 2024; Bostrom & Durrett, 2020; Sun et al., 2023). For instance, Allen-Zhu & Li (2025) observe that models may be aware of their mistakes without being able to correct them, while Allen-Zhu & Li (2024) highlight how subtle data modeling choices can impede fine-tuning progress. Collectively, these insights suggest that models often possess the latent capacities needed to tackle ARC-AGI; the real challenge is creating the conditions under which these capacities can be reliably expressed.
@@
In the sections that follow, we detail our data modeling and training strategies, describe our DFS-based solution exploration, and provide comprehensive results and ablation studies.
Our final model, along with the training and inference code, is publicly available on GitHub.
@@
Ryan Greenblatt (Greenblatt, 2024) 42
Jeremy Berman (Berman, 2024) 58.5
@@ 2 Related Work
The Abstraction and Reasoning Corpus (ARC) has played a central role in advancing research on abstract reasoning in artificial intelligence, inspiring a wide range of studies focused on its dataset, competitive benchmarks, and the development of solutions driven by resource constraints.
@@
The Abstraction and Reasoning Corpus (ARC-AGI) introduced by Chollet (2019) challenges the idea that language models can efficiently generalize from a small number of examples, often referred to as few-shot prompting. The original ARC-AGI dataset consists of 900 reasoning tasks, divided into 400 training tasks, 400 public evaluation tasks, and 100 private and thus unpublished evaluation tasks. Each task involves input and output grids of varying sizes, ranging from 1x1 to 30x30 and utilize a palette of ten distinct colors.
The objective of each individual ARC task is to discern the transformation rule from input to output from the examples and apply it to new input grids to generate the correct outputs. A task is considered successfully solved when the model produces the accurate output within a maximum of two attempts. Designed to be straightforward for humans yet challenging for machine learning systems, the tasks highlight the current limitations of AI in abstract reasoning. In a study by LeGris et al. (2024), the average human was able to correctly solve 60.2% of the evaluation tasks, while 97.8% of the tasks were solved by at least one participant using two guesses.
Figure 2: Our standard tokenization approach. Note that we use one token per cell instead of compressing the problem more. We also try to not include any unnecessary delimiters. The Pre-prompt (the alphabet in upper then lower case, i.e. “A…Za…z”) is only included for the first example. Depending on the model and run there might be some small changes to the pre-prompt and input/output prefix tokens.
@@
Other approaches explored various strategies for utilizing LLMs. In Li et al. (2025), the authors classify two different avenues: Induction, where a LLM infers a function that can solve the problem which is then applied (often using python or a DSL), and Transduction, where the LLM directly generates the solution using a tokenized description of the problem (see Figure 2). The authors argue that these approaches solve different kinds of problems, despite using the same underlying architecture. In their experiments, induction and transduction solve roughly the same amount of problems ($38\%$ and $43\%$ respectively), which can be increased to $56.75\%$ by employing ensembles. Additionally, they use the induction network to generate a large set of novel challenges, dubbed ARC-Heavy. Some approaches make use of alternative ARC datasets, such as ConceptARC (Moskvichev et al., 2023). The most notable - and only additional dataset we use - is the well-known RE-ARC dataset. Hodel (2024) introduced this dataset, which implements generators for all $400$ tasks of the public training dataset. Their code can be used to produce an arbitrary amount of training data for these tasks, but does not introduce novel challenges. All other datasets might include challenges that mimic the evaluation challenges - thereby reducing the difficulty of those challenges immensely. By only using the RE-ARC dataset, we still increase our training data immensely, but stay close to solving the ARC challenge as intended.
Data augmentation has been a common approach in previous ARC-AGI competitions (Akyürek et al., 2024; Li et al., 2025). However, our method extends beyond traditional dataset augmentation, applying transformations throughout our approach, during training (initial finetuning as well as test time training), inference and selection.
@@ 3 Notations and Setup
over candidate solutions $s\in\mathcal{S}_{p}$. If exactly one possible valid answer exists the distribution $P(\cdot\mid p)$ would be sharply peaked at $s^{*}_{p}$. While this is the case for most challenges, we will keep our theory more general, assuming multiple valid answers might exist. This can also arise from insufficient information in the given example pairs, which in the worst case prevents us from uniquely inferring the correct solution based solely on the provided data. Examples for this are sometimes found in ARC-AGI, which frequently results in an update of the dataset (Neoneye, 2024; RubenKelevra, 2024).
@@
Figure 4: Top-2 accuracy and coverage of different selection methods as a function of the confidence threshold $T$. Solid colored lines denote the fraction of tasks solved using a specific selection algorithm. The solid black line shows the fraction of tasks where the correct solution was among the sampled candidates, and thereby provides an upper bound for the performance of the selection algorithms. The dotted lines evaluate the performance of our selection algorithms, compared to this upper bound: What percentage of correct candidates are actually selected when they are present? It shows that even when using low DFS probabilities - and therefore sampling a high number of candidates - PoE is able to select the correct solution among all candidates with high specificity.
@@ 4 Methods
The trained LLM then provides us with a probability distribution $\hat{P}(\cdot\mid p)$ over solutions $\mathcal{S}_{p}$. Using multinomial sampling, this allows us to sample $s\sim\hat{P}(\cdot\mid p)$ However, sampling repeatedly from $\hat{P}$ may be expensive and does not ensure coverage of high-probability solutions. In contrast, enumerating $\mathcal{S}_{p}$ in full would provide us with full knowledge of $\hat{P}(s)$ but is impractical. Instead, we rely on a more systematic procedure to select promising candidates by deriving a candidate set of high-probability solutions via threshold-based search. Subsequently, we refine their probabilities by aggregating over multiple problem augmentations.
@@
However, once we have generated the set $\mathcal{C}_{p,T}$, the highest-probability solution according to a single augmentation is not always correct.
This limitation is partially caused by the autoregressive architecture, as models can only attend to previously generated tokens when predicting the next token. This constraint means optimal decisions sometimes require information that becomes available only in later predictions. In the Sudoku experiment (Section 5.6), for instance, the model may need to solve the entire puzzle internally before predicting the first cell, potentially leading to confident but incorrect early predictions. Once an error occurs, the model cannot recover since subsequent predictions build on the incorrect foundation. The model lacks training for stability under prediction errors, causing cascading mistakes with unexpectedly high confidence scores. This explains why the highest probability sequence from a single forward pass may not correspond to the globally optimal solution.
@@
This two-step approach; (1) DFS-based generation with single-augmentation pruning and (2) post-hoc multi-augmentation scoring, ensures that we systematically explore high-probability solutions and then refine their rankings, accounting for LLM inconsistencies across problem representations. In practice, even if several solutions enter $\mathcal{C}_{p,T}$, their final ranks can vary greatly. By consolidating evidence from multiple perspectives, the correct solution often stands out and becomes easier to pinpoint.
@@ 4.1 Product of Expert Augmentations
See Appendix C for a proof of Theorem 4.1. The key takeaway is that $\log Z\leq 0$ becomes smaller whenever augmentations disagree, which can improve the ensemble in expectation relative to any random single-augmentation predictor. As a result, this approach performs especially well when different experts disagree - a state which naturally arises in our case, due to the causal autoregressive nature of the LLMs.
@@
Table 2: Two-guess-accuracy on the ARC-AGI public evaluation set when adding parts of our method. Baseline score shows performance of our network after initial fine-tuning, generating two samples with stochastic sampling. TTT adds test-time training. 16xAug samples one solution candidate for each of 16 random augmentations of each task, choosing the two with highest sampling probability as guesses. PoE uses the product of experts to select the two best of the 16 sampled candidates, again using 16 (different) random augmentations to calculate the PoE score. Finally, DFS leverages our custom depth-first-search sampling scheme with $T=9\%$ for candidate generation.
@@ 5.1 Data Modeling
First, due to the limited context size in typical LLM architectures, an increase of inference time and decline in performance on long context tasks (Liu et al., 2024), we require a representation that minimizes the number of tokens the model needs to process. Secondly, it is widely recognized that numerous common failure modes in Large Language Models (LLMs) stem from tokenization (Singh & Strouse, 2024; Bostrom & Durrett, 2020; Sun et al., 2023). For instance, standard tokenization techniques group numbers (some but not all combinations) of one, two or three succeeding digits into dedicated “grouped-digit tokens” (Singh & Strouse, 2024). These kinds of merges would complicate the puzzles unnecessarily.
@@
As illustrated in Figure 2, we add a small number of extra tokens to the start of a task. Surprisingly, this addition slightly improves the model’s performance. We believe that during fine-tuning (where the embedding layers are also trained), the model learns to use these extra tokens as a form of computational buffer, which influences every subsequent token, thereby enhancing overall performance.
@@ 5.2 Training the models
Therefore, we used Low-Rank Adaptation (LoRA) (Hu et al., 2022), 4-bit quantization and gradient checkpointing, all supported by the unsloth library. We applied the LoRA adaptations to all layers of the network, including the input and output embeddings.
@@
Initial fine-tuning: The initial fine-tuning used a LoRA rank of $256$ and was done on a single H100 GPU. While several ARC-like datasets exist, such as ConceptARC (Moskvichev et al., 2023) and ARC-Heavy (Li et al., 2025), we elect to only use RE-ARC (Hodel, 2024) for training. This is done to minimize ”conceptual leakage”, where a particular type of problem might be present in the training data, reducing the difficulty of the evaluation tasks substantially in a way that was not intended. Instead, we train only on replications of the training examples of the offical ARC-AGI training set (i.e. RE-ARC), minimizing this effect and making sure that our results are robust.
@@
The initial fine-tuning took 98 GPU hours on a Nvidia H100, while test-time training takes (on average) 51 seconds for a single task on a Nvidia RTX 4090 GPU. For an overview of our training parameters, see Table 4 in the Appendix.
@@ 5.3 Solution Inference
As introduced in Section 4, we generate potential solution candidates using DFS-based sampling to produce the set $C_{p,T}$. The goal here is to generate a small set of candidates with a high chance of containing the correct solution - and doing so quickly. Our set of augmentations $\Phi$ includes $16$ functions per task - each $D_{8}$ symmetry is used twice but with different, randomly chosen color permutations and example re-orderings. Note that this is the same class of augmentations as used in training, but each color permutation and example ordering is newly randomized. Table 3 provides a comparison between DFS, Beam-search, multinomial and greedy sampling. DFS sampling is able to quickly and efficiently find a high quality set of candidates, while having low computational overhead compared to stochastic sampling for generating multiple solutions and using substantially less VRAM than beam search. In addition, it exhibits a lower false positive rate. While DFS with $T=9\%$ finds less correct solutions than 4x Stochastic sampling ($76.0\%$ vs $77.3\%$), it still results in a better selection score, as it, on average, only returns about half as many false positives. Moreover, DFS accomplishes this using only a fourth of the inference time (9:32h vs 39:47h).
Table 3: Comparison of sampling and selection strategies on the 400 tasks of the ARC-AGI public evaluation set: Under “Candidate generation”, we list the percentage of correct solutions sampled with different strategies using 16 augmented versions (reflections, rotations, and randomly permuted colors and examples) of each task. We also list the average number of candidates generated per task, the runtime of the sampling process on the full dataset and the maximum video memory consumption. Under “Selection”, we compare the accuracy of various selections strategies, performed on the scores calculated in a subsequent scoring process on 16 additional random augmentations. Total runtime includes the test-time fine-tuning on a task’s examples (see Table 4 in the appendix), as well as the candidate generation and selection process. All experiments were performed with the NeMo-Minitron-8B model on a Nvidia RTX 4090 GPU.
@@
Comparison to beam search: While beam search with 4 beams can achieve the same accuracy as DFS with $T=9\%$, it requires roughly twice the amount of VRAM (7.3GB vs 14GB), as it explores four paths simultaneously, while DFS only needs to keep a single path in memory at any time. It also takes four times as long (37:36h vs 9:32h) for the candidate generation step. The speed advantage of DFS comes mostly from early pruning of low probabilty paths. In Beam Search, the same amount of paths is explored each time, regardless of their cumulative sampling probability, while DFS stops when the cumulative sampling probability falls beyond the chosen threshold, thereby reducing unnecessary computations. Additionally, for all augmentations after the first one, we pass the most promising solution candidate found so far as an inital guess to the DFS and process it in a single forward pass before starting backtracking, which is much faster than token-by-token generation. Note that these comparisons should be interpreted with caution, as the beam search algorithm is not implemented in the unsloth library used for the other experiments, which might provide some time savings. However, even when accounting for those savings, beam search still requires far more time overall, as it returns a significant amount of false positives, which increase the runtime required in the subsequent selection process, where each candidate is evaluated under different augmentations.
We compare the number of candidate sets that contain the correct solution for different values of $T$ in Figure 4. This function is monotonically increasing in $T$, but so are inference costs and the size of the set $C_{p,T}$, making the selection of the correct candidate harder. Our final results are calculated using $T=9\%$, as it uses roughly the same amount of inference time as greedy sampling.
We provide pseudo-code for our DFS sampling algorithm in Algorithm 1 in the Appendix.